Could someone please help me solve this problem?
A metal stays solid at Fahrenheit temperature below 1865.9 degrees. Determine ( in terms of an inequality) those Celsius temperature for which metal stays solid, use formula F=9/5c+32
the solution set is {clc<_ degrees}, round to nearest tenth

Answer by Alan3354(69443)

C = (5/9)*(F - 32)
C = (5/9)*(1865.9 - 32) = 5*1833.9/9
C = 1018.8333...
Solid for temperature < 1018.833 degs C